OpenStudy (karatechopper):

Moving lava can build up and form beaches at the coast of an island. The growth of an island in a seaward direction may be modeled as 8y+2 centimeters, where y represents the number of years that the lava flows. An island has expanded 60 centimeters seaward. How long has the lava flowed?

OpenStudy (karatechopper):

ur here... so help me?!

OpenStudy (ybarrap):

growth = g = 8y +2 cm y = number of years if g = 60 cm, what is ynew? Where ynew is how long it took to expand 60 cm So, g = 60 = 8ynew + 2 8ynew + 2 = 60 8ynew = 60 - 2 ynew = 58/8 = 29/4 years Ans. 29/4 years, exactly.

OpenStudy (karatechopper):

thaks